# Investigating the Electroweak Phase Transition with a Real Scalar Singlet at a Muon Collider

### v1

@@ 2 Electroweak Phase Transition with a Scalar Singlet
Where ${\left(\frac{\sqrt{C}}{\mu_{h}}\right)^{SM}\approx 6.6\times 10^{-3}}$. For large negative values for $\alpha$, $\alpha>\frac{1}{2}\lambda_{hs}\omega$, the square bracket in Eq.(28) becomes greater than one; however $\mu_{h}$ simultaneously decreases and the $C$-term increases, which collectively leads to a significant decrease in the critical temperature in xSM, $T_{c}^{\text{\tiny xSM}}<T_{c}^{\text{\tiny SM}}$.
This approximate analytical investigation of the xSM extension is corroborated by the numerical methods employed to account for the full NLO (CW and finite temperature) corrections to the TL potential at different renormalization scales (conservative and liberal categories) in order to address the theoretical uncertainties in the FOEWPT parameter space points, which are necessary for accurately exploring the muon collider. As illustrated in Fig.2, the improvement of the sphaleron condition by the xSM extension is evident, as is the decrease in critical temperature due to the singlet contribution to the $C$ and $\mu_{h}^{2}$-terms, as demonstrated in Fig.3 for various singlet mass values.
@@ 2.1 First-Order Electroweak Phase Transition Parameter Space within the xSM
Catalyzing a first-order EWPT through the real singlet extension depends on the parameter space points, particularly on the $\mathbb{Z}_{2}$-asymmetric parameters and Higgs-scalar quartic coupling, as shown in Eqs. (26). The additional scalar degree of freedom introduces five free parameter $\{\mu_{s}^{2},\,\lambda_{s},\,\alpha,\,\beta,\,\lambda_{hs}\}$ which need to be determined. In profumo2007singlet ; carena2020electroweak ; espinosa2012strong , a set of diverse constraints, stemming from theoretical and phenomenological sources, were discussed to restrict these parameters,
@@
  • The validity of perturbative expansion requires quartic couplings to be smaller than unity, $(\lambda_{h},\lambda_{s},\lambda_{hs})\in[0:1]$.
  • Electroweak symmetry breaking necessitates a positive determinant of the mass matrix in Eq. (9) as ($v_{\text{\tiny EW}},\omega_{\text{\tiny EW}}$) should constitute the true global minima, which requires, $04\lambda_{h}\lambda_{s}-\lambda_{hs}>$ 0.
@@
We used the values obtained in  Papaefstathiou_2021 , in which nearly identical constraints were applied with minor variations. In this study, valid points satisfying the sphaleron condition $\frac{v_{c}}{T_{c}}>1$ were obtained through a scan using the PhaseTracer package athron2020phasetracer , which tracks vacuum evolution by evaluating the thermal integral of Eq. (11) numerically by employing the methods described in  fowlie2018fast ; wainwright2012cosmotransitions , rather than the HTE approximation discussed in the previous calculations to obtain an approximate understanding of xSM. Furthermore, the ranges of the “portal” coupling $\lambda_{hs}$, and singlet quartic coupling $\lambda_{s}$, were set to be quite loose in Papaefstathiou_2021 , as listed in Table 2. The rationale behind this approach is that for colliders to exclude the possibility of a strong first-order EWP, the candidate parameter space points should be relatively broad.
@@
The values of the $\lambda_{hs}$ coupling control the shift in the quadratic correction on $C$ from the SM expectation, and consequently improve the sphaleron condition. However, this raises concerns regarding the validity of the perturbative expansion and its sensitivity to theoretical uncertainties related to the scale and gauge dependencies of the thermal parameters. This consideration is crucial not only for analyzing specific points but also for comprehending the parameter space lewicki2024impacttheoreticaluncertaintiesmodel ; gould2023higherorderscosmologicalphase . According to Papaefstathiou_2021 , the dominant uncertainty originates from the slow convergence of perturbation theory which is manifested in the dependence on an arbitrary renormalization scale. We follow Papaefstathiou_2021 and define a conservative point as a point that admits an SFOEWPT for a full order of magnitude variation of the RG scale, whereas a liberal point is one that admits an SFOEWPT for particular choices of the scale [Check Papaefstathiou_2021 for more details].
@@
A substantial number of parameter space points do fulfill the SFOEWPT (sphaleron) condition out of the $\mathcal{O}(10^{6})$ points generated according to Table 2, as obtained by PhaseTracer. The SFOEWPT condition is plotted against the scalar mass in Fig. 2 for conservative and liberal categories. The observation that conservative category values shift more towards lower values of scalar mass, unlike liberal values, is a good indicator that this category can be verified by current collider searches. For instance, the LHC searches for an SM-like Higgs boson cms2013search , which may lead to the exclusion of the conservative category.
@@
Portal coupling $\lambda_{hs}$ improves the SFOEWPT condition primarily through the correction it contributes to the quadratic $C$-term, which subsequently reduces the critical temperature. This phenomena is illustrated in Fig. 3 (left panel) which demonstartes the increase in the $C-$term with $\lambda_{hs}$, contrary to the SM expectation. The right panel of Fig. 3 depicts the tendency towards a lower critical temperature compared to the SM expectation, for both Conservative and Liberal categories. Furthermore, the valid parameter points exhibit a propensity to concentrate towards lower values of the quartic couplings ($\lambda_{s},\,\lambda_{hs}$), as shown in Fig. 4, which indicates favorable behavior regarding the validity of perturbative expansion.
@@ 3 The Electroweak Phase Transition at Colliders
A direct test for SFOEWPT could be derived from the gravitational waves emanating from bubble dynamics. Specifically, sound waves in the plasma and magneto-hydrodynamic turbulence generate gravitational waves that are expected to peak at low frequencies ranging from mHz to 10 Hz. Collider searches provide another promising and complementary approach for verifying the nature of EWPT. This is attributable to the fact that the catalyst of the FOEWPT, the singlet scalar in this instance, can be probed either directly through resonant production, if its mass is within the range of current and future searches, or indirectly, through precise measurements of potential deviations in the Higgs boson couplings from the SM expectation. In contrast to gravitational wave detection, which can only be produced from FOEWPT, collider searches can be sensitive to SOEWPT as well; however, this is more challenging owing to the weak mixing with the Standard Model particles.
@@ 3.1 Direct Singlet Production at Colliders
The mass of the new scalar particle in xSM is expected to be of $\mathcal{O}(T_{\text{\tiny EW}})$, which suggests the possibility of resonant production at colliders ramsey2020electroweak . At zero temperature, that is, in the collider scenario, the mass of the singlet is given by $m_{2}(v_{\text{\tiny EW}},\omega_{\text{\tiny EW}})$, Eq. (10), which yields a scalar mass in the range $m_{s}\in[200,1000]$, when utilizing the range of parameters as shown in Fig. 2. This can be approximated using the curvature of the effective potential in Eq. (21),
@@
Because of the anticipated low mass of the new scalar, $\lesssim\mathcal{O}$(TeV), both hadron and lepton colliders are valid options for detection. Previous studies papaefstathiou2022electro ; ramsey2020electroweak have confirmed that a 100 TeV proton-proton collider can substantially explore this parameter space. Recently, there has been a growing interest in muon colliders, which can explore the same parameter space efficiently, at a significantly lower center of mass energy ($\sqrt{s}\in[3-10]$ TeV) ramsey2020electroweak ; liu2021probing . In a muon collider, most of the available energy is consumed in the hard process, which directly relates the resonance peak to the center of mass energy according to $\frac{\sqrt{s}}{2m_{s}}\approx 1.7$ as demonstrated in ref. ramsey2020electroweak . This characteristic is unique for lepton colliders, unlike hadron colliders where $E_{\text{\tiny CM}}$ of the beam differs from that of the parton beams, necessitating integration over the parton density functions. Furthermore, a muon collider will not suffer from the huge QCD radiation resulting from the initial state as in the case of hadron colliders, nor will its energy be dissipated in the large synchrotron radiation as in an electron collider. Notably, the optimal channel for the single production of a new scalar is through vector boson fusion (VBF), mainly $W^{+}W^{-}$ fusion, which contributes $\sim 90\%$ of the total cross section liu2021probing ; forslund2022high .
@@ 3.2 Indirect Evidence: Modification of Higgs Boson Couplings
which consequently form a “portal” of the new scalar to the other SM particles. The other self-interaction couplings of the new scalar do play a crucial role in determining the points of EWPT but do not substantially affect phenomenological studies  [^3] For example, the $\lambda_{s}$ coupling only appears in the daisy resummation part of the effective potential curtin2014testing .. The terms in Eq. (30) inevitably lead to mixing between $H$ and $S$ and consequently the new scalar “inherits” interactions with SM particles. The mass eigenstates are obtained using the rotation angle $\theta$, after symmetry breaking as follows:
@@
$\tan\theta=\frac{M^{2}_{hs}}{M^{2}_{h}-M^{2}_{s}+\sqrt{(M^{2}_{h}-M^{2}_{s})^{2}+M^{2}_{hs}}}\;,$ (32)
where $M^{2}_{h},\,M^{2}_{s},\,M^{2}_{hs}$ are the components of the mass matrix in Eq. (9). Consequently, the SM-like Higgs boson couplings must be adjusted according to:
@@
Here, $g_{\text{\tiny$h_{1}$XX}}^{\text{\tiny SM}}\cos\theta$ represents the scaling of the original Higgs boson couplings, and -$g_{\text{\tiny$h_{2}$XX}}^{\text{\tiny SM}}\sin\theta$ represents the new scalar couplings to the $XX$-SM-particles. This opens up a range of possible precision tests such as small deviations in SM-like Higgs boson production rates according to $\sigma^{\text{\tiny BSM}}_{h}=\cos^{2}\theta\sigma^{\text{\tiny SM}}_{h_{1}}$. Another significant phenomenological test would be the deviation of Higgs boson trilinear self-coupling from the SM expectation, where after symmetry breaking this term becomes,
@@ 4 Constraints on the xSM at a Muon Collider
The dominant production mechanism of the new scalar $(S)$ at a muon collider is vector boson fusion (VBF), with $W^{+}W^{-}$ fusion (WWF) dominating $ZZ$ fusion (ZZF) as reported in previous studies forslund2022high ; liu2021probing ; costantini2020vector .
Figure 6: The different production mechanisms are illustrated in (A), which also shows that VBF is dominated by $W^{+}W^{-}$ fusion (WWF) (green curve), which is further confirmed in (B), showing that WWF accounts nearly for 90% of VBF. (B) also shows that the $S$-production cross section is only marginally sensitive on its mass $M_{S}$. The plots have been generated through MadGraph5_aMC@NLO simulations at parton level.
@@ 4.1 Event Generation
Both signal and background Monte Carlo events were generated at the parton level using MadGraph5_aMC@NLO Alwall:2011uj (MG5_aMC), while the scalar decays to gauge/Higgs bosons were generated in HERWIG 7 Bahr:2008pv ; Bellm:2017bvx ; Gieseke:2011na ; Arnold:2012fq ; Bellm:2013hwb ; Bellm:2019zci ; Bewick:2023tfi , along with the parton showering, underlying event and hadronization. The HwSim plugin for HERWIG 7 hwsim was used to generate ROOT Brun:1997pa files for all event samples. Signal events were produced in MG5_aMC by using the loop_sm_scalar model Papaefstathiou_2021 . [^4] We used electron-positron beams for event generation instead of muon-antimuon beams since HERWIG 7, at the time of writing, cannot recognize muons in the initial state. This does not affect our analysis at all due to the lepton universality and the irrelevance of the lepton mass at high energies. Due to background events at a muon collider occurring at much lower rates than at a hadron collider, we took into account all the possible background events for each final state by generating it directly in MG5_aMC, without importing the loop_sm_scalar model. For example, for $\mu^{+}\mu^{-}\rightarrow\mu^{+}\mu^{-}S\rightarrow\mu^{+}\mu^{-}l^{+}l^{+}l^{-}l^{-}$, we generated the combined background, $\mu^{+}\mu^{-}\rightarrow\mu^{+}\mu^{-}l^{+}l^{+}l^{-}l$.
@@
Combined process generation implicitly contains all the dominant backgrounds resulting from $ZZ,\,Z\tau^{+}\tau^{-},\,Zt\tilde{t},\cdots$. We have examined this claim by comparing the kinematic distributions ($P_{T},\,\eta$) of the $4l$-combined background to the dominant $4l$ background sources, such as those coming from $ZZ,\,Z\tau^{+}\tau^{-},\,Zt\tilde{t}$, and confirmed that the combined $4l$- background is indeed almost identical to the sum of the individual dominant backgrounds, as evident in Fig. 7. Therefore, in our analysis, we compare the signals to the combined backgrounds, instead of just the main ones.

### v2

@@ 2 Electroweak Phase Transition with a Scalar Singlet
Where ${\left(\frac{\sqrt{C}}{\mu_{h}}\right)^{SM}\approx 6.6\times 10^{-3}}$. For large negative values for $\alpha$, |$\absolutevalue{\alpha}>\frac{1}{2}\lambda_{hs}\omega$, the square bracket in Eq. (28) becomes greater than one; however $\mu_{h}$ simultaneously decreases and the $C$-term increases, which collectively leads to a significant decrease in the critical temperature in xSM, $T_{c}^{\text{\tiny xSM}}<T_{c}^{\text{\tiny SM}}$.
This approximate analytical investigation of the xSM extension is corroborated by the numerical methods employed to account for the full NLO (CW and finite temperature) corrections to the tree-level potential at different renormalization scales (conservative and liberal categories, see later) in order to address the theoretical uncertainties in the FOEWPT parameter space points, which are necessary for accurately exploring the muon collider. As illustrated in Fig. 2, the improvement of the sphaleron condition by the xSM extension is evident, as is the decrease in critical temperature due to the singlet contribution to the $C$ and $\mu_{h}^{2}$-terms, as demonstrated in Fig. 3 for various singlet mass values.
@@ 2.1 First-Order Electroweak Phase Transition Parameter Space within the xSM
Catalyzing a FOEWPT through the real singlet extension depends on the parameter space points, particularly on the $\mathbb{Z}_{2}$-asymmetric parameters and the Higgs-scalar quartic coupling, as shown in Eqs. (26). The additional scalar degree of freedom introduces five free parameter $\{\mu_{s}^{2},\,\lambda_{s},\,\alpha,\,\beta,\,\lambda_{hs}\}$ which need to be determined. In profumo2007singlet ; carena2020electroweak ; espinosa2012strong ; fernandez2023nu , a set of diverse constraints, stemming from theoretical and phenomenological sources, were discussed to restrict these parameters,
@@
  • The validity of perturbative expansion requires quartic couplings to be smaller than unity, $(\lambda_{s},\lambda_{hs})\in[0:4\pi]$.
  • Electroweak symmetry breaking necessitates a positive determinant of the mass matrix in Eq. (9) as ($v_{\text{\tiny EW}},\omega_{\text{\tiny EW}}$) should constitute the true global minimum, which requires, $04\lambda_{h}\lambda_{s}-\lambda_{hs}>$ 0.
@@
We used the values obtained in Ref. Papaefstathiou_2021 , in which nearly identical constraints were applied with minor variations. In this study, valid points satisfying the sphaleron condition $\frac{v_{c}}{T_{c}}>1$ were obtained through a scan using the PhaseTracer package athron2020phasetracer , which tracks vacuum evolution by evaluating the thermal integral of Eq. (11) numerically by employing the methods described in  fowlie2018fast ; wainwright2012cosmotransitions , rather than the HTE approximation discussed in the previous calculations to obtain an approximate understanding of xSM. Furthermore, the ranges of the “portal” coupling $\lambda_{hs}$, and singlet quartic coupling $\lambda_{s}$, were set to be quite loose in Papaefstathiou_2021 , as listed in Table 2. The rationale behind this approach is that for colliders to exclude the possibility of a strong FOEWPT (SFOEWPT), the candidate parameter space points should be relatively broad.
@@
The values of the $\lambda_{hs}$ coupling control the shift in the quadratic correction on $C$ from the SM expectation, and consequently improve the sphaleron condition. However, this raises concerns regarding the validity of the perturbative expansion and its sensitivity to theoretical uncertainties related to the scale and gauge dependencies of the thermal parameters. This consideration is crucial not only for analyzing specific points, but also for comprehending the parameter space lewicki2024impacttheoreticaluncertaintiesmodel ; gould2023higherorderscosmologicalphase . According to Papaefstathiou_2021 , the dominant uncertainty originates from the slow convergence of perturbation theory which is manifested in the dependence on an arbitrary renormalization scale. We follow Ref. Papaefstathiou_2021 and define a “conservative” point as a point that admits an SFOEWPT for a group of eight scale and gauge variations, with at least one of these variations possessing a Higgs field vev within $246\pm 30$ GeV, that also corresponds to a transition to the absolute (i.e. deepest) minimum of the potential. On the other hand a “liberal” point satisfies SFOEWPT for one of the eight scale and gauge variations and the Higgs vev condition for any other. See Ref. Papaefstathiou_2021 for further details.
@@
A substantial number of parameter space points fulfill the SFOEWPT sphaleron condition out of the $\mathcal{O}(10^{6})$ points generated according to Table 2. The SFOEWPT condition is plotted against the scalar mass in Fig. 2 for conservative and liberal categories. The observation that conservative category values shift more towards lower values of scalar mass, unlike liberal values, is a good indicator that this category can be verified by current collider searches. For instance, the LHC searches for an SM-like Higgs boson cms2013search , which may lead to the exclusion of the conservative category.
@@
The portal coupling $\lambda_{hs}$ improves the SFOEWPT condition primarily through the correction it contributes to in the quadratic $C$ term, which subsequently reduces the critical temperature. This phenomenon is illustrated in Fig. 3 (left panel) which demonstrates the increase in the $C-$term with $\lambda_{hs}$, contrary to the SM expectation. The right panel of Fig. 3 shows the tendency toward a lower critical temperature compared to the SM expectation, for both the Conservative and Liberal categories. Furthermore, the valid parameter points exhibit a propensity to concentrate towards lower values of the quartic couplings ($\lambda_{s},\,\lambda_{hs}$), as shown in Fig. 4, which indicates favorable behavior regarding the validity of perturbative expansion.
@@ 3 The Electroweak Phase Transition at Colliders
A direct test for SFOEWPT could be derived from the gravitational waves emanating from bubble dynamics. Specifically, sound waves in the plasma and magneto-hydrodynamic turbulence generate gravitational waves that are expected to peak at low frequencies ranging from mHz to 10 Hz. Collider searches provide another promising and complementary approach for verifying the nature of EWPT. This is attributable to the fact that the catalyst of the FOEWPT, the singlet scalar field in this instance, can be probed either directly through resonant production of a new scalar, if its mass is within the range of current and future searches, or indirectly, through precise measurements of potential deviations in the Higgs boson couplings from the SM expectation. In contrast to gravitational wave detection, which can only be produced from FOEWPT, collider searches can be sensitive to SFOEWPT as well; however, this is more challenging owing to the weak mixing with the Standard Model particles.
@@ 3.1 Direct Singlet Production at Colliders
The mass of the new scalar particle in xSM is expected to be of $\mathcal{O}(T_{\text{\tiny EW}})$, which suggests the possibility of resonant production at colliders ramsey2020electroweak . At zero temperature, i.e. in the collider scenario, the mass of the singlet-like scalar is given by $m_{2}(v_{\text{\tiny EW}},\omega_{\text{\tiny EW}})$, Eq. (10), which yields a scalar mass in the range $m_{s}\in[200,1000]$ GeV, when utilizing the range of parameters as shown in Fig. 2. This can be approximated using the curvature of the effective potential in Eq. (21),
@@
Due to the anticipated low mass of the new scalar, $\lesssim\mathcal{O}$(TeV), both hadron and lepton colliders are valid options for detection. Previous studies papaefstathiou2022electro ; ramsey2020electroweak have confirmed that a 100 TeV proton-proton collider can substantially explore this parameter space. Recently, there has been a growing interest in muon colliders, which can explore the same parameter space efficiently, at a significantly lower center of mass energy ($\sqrt{s}\in[3-10]$ TeV) ramsey2020electroweak ; liu2021probing . In a muon collider, most of the available energy is consumed in the hard process, which directly relates the resonance peak to the center of mass energy according to $\frac{\sqrt{s}}{2m_{s}}\approx 1.7$ as demonstrated in ref. ramsey2020electroweak . This characteristic is unique for lepton colliders, unlike hadron colliders where $E_{\text{\tiny CM}}$ of the beam differs from that of the parton beams, necessitating integration over the parton density functions. Furthermore, a muon collider will not suffer from the huge QCD radiation resulting from the initial state as in the case of hadron colliders, nor will its energy be dissipated in the large synchrotron radiation as in an electron collider. Notably, the optimal channel for the single production of a new scalar is through vector boson fusion (VBF), mainly $W^{+}W^{-}$ fusion, which contributes $\sim 90\%$ of the total cross section liu2021probing ; forslund2022high .
@@ 3.2 Indirect Evidence: Modification of Higgs Boson Couplings
which consequently form a “portal” of the new scalar to the other SM particles. The other self-interaction couplings of the new scalar do play a crucial role in determining the points of EWPT but do not substantially affect phenomenological studies. [^3] For example, the $\lambda_{s}$ coupling only appears in the daisy resummation part of the effective potential curtin2014testing . The terms in Eq. (30) inevitably lead to mixing between $H$ and $S$ and consequently the new scalar “inherits” interactions with SM particles. The mass eigenstates are obtained using the rotation angle $\theta$, after symmetry breaking as follows:
@@
$\cot\theta=\frac{2M^{2}_{hs}}{M^{2}_{hh}-M^{2}_{ss}+\sqrt{(M^{2}_{hh}-M^{2}_{ss})^{2}+4M^{4}_{hs}}}\;,$ (32)
where $M^{2}_{hh},\,M^{2}_{ss},\,M^{2}_{hs}$ are the components of the mass matrix in Eq. (9). Consequently, the SM-like Higgs boson couplings must be adjusted according to:
@@
Here, $g_{\text{\tiny$h_{1}$XX}}^{\text{\tiny SM}}\cos\theta$ represents the scaling of the original Higgs boson couplings, and $g_{\text{\tiny$h_{2}$XX}}^{\text{\tiny SM}}\sin\theta$ represents the new scalar couplings to the $XX$-SM-particles. This opens up a range of possible precision tests such as small deviations in SM-like Higgs boson production rates according to $\sigma^{\text{\tiny BSM}}_{h}=\cos^{2}\theta\times\sigma^{\text{\tiny SM}}_{h_{1}}$. Another significant phenomenological test would be the deviation of Higgs boson trilinear self-coupling from the SM expectation, where after symmetry breaking this term becomes,
@@ 4 Constraints on the xSM at a Muon Collider
The dominant production mechanism of the new scalar $(S)$ at a muon collider is vector boson fusion (VBF), with $W^{+}W^{-}$ fusion (WWF) dominating $ZZ$ fusion (ZZF) as reported in previous studies buttazzo2018fusing ; forslund2022high ; liu2021probing ; costantini2020vector .
Figure 6: The different production mechanisms are illustrated in (A), which also shows that VBF is dominated by $W^{+}W^{-}$ fusion (WWF) (green curve), which is further confirmed in (B), showing that WWF accounts nearly for 90% of VBF. (B) also shows that the $S$-production cross section is only marginally sensitive on its mass $M_{S}$. The plots have been generated through MadGraph5_aMC@NLO simulations at parton level for $\sin\theta=1$.
@@ 4.1 Event Generation
Both signal and background Monte Carlo events were generated at the parton level using MadGraph5_aMC@NLO Alwall:2011uj (MG5_aMC), while the scalar decays to gauge/Higgs bosons were generated in HERWIG 7 Bahr:2008pv ; Bellm:2017bvx ; Gieseke:2011na ; Arnold:2012fq ; Bellm:2013hwb ; Bellm:2019zci ; Bewick:2023tfi , along with the parton showering, underlying event and hadronization. The HwSim plugin for HERWIG 7 hwsim was used to generate ROOT Brun:1997pa files for all event samples. Signal events were produced in MG5_aMC by using the loop_sm_scalar model Papaefstathiou_2021 . [^4] We used electron-positron beams for event generation instead of muon-antimuon beams since HERWIG 7, at the time of writing, cannot readily handle muons in the initial state. This does not affect our analysis at all because of lepton universality, and the irrelevance of the lepton mass at high energies.
@@
Due to background events at a muon collider occurring at much lower rates than at a hadron collider, we took into account all the possible background events for each final state by generating it directly in MG5_aMC, without importing the loop_sm_scalar model. For example, for $\mu^{+}\mu^{-}\rightarrow\mu^{+}\mu^{-}S\rightarrow\mu^{+}\mu^{-}l^{+}l^{+}l^{-}l^{-}$, we generated the combined background, $\mu^{+}\mu^{-}\rightarrow\mu^{+}\mu^{-}l^{+}l^{+}l^{-}l$. Combined process generation implicitly contains all the dominant backgrounds resulting from $ZZ,\,Z\tau^{+}\tau^{-},\,Zt\tilde{t},\cdots$. We have examined this claim by comparing the kinematic distributions ($P_{T},\,\eta$) of the $4l$-combined background to the dominant $4l$ background sources, such as those coming from $ZZ,\,Z\tau^{+}\tau^{-},\,Zt\tilde{t}$, and confirmed that the combined $4l$- background is indeed almost identical to the sum of the individual dominant backgrounds, as is evident in Fig. 7. Therefore, in our analysis, we compare the signals to the combined backgrounds, instead of just the main ones.